# Ruishu Liu, Yulin Cao,  Xiaojie Wang

### v1

@@ 3 The explicit positivity-preserving Euler-type scheme
$\mathbb{E}\Big{[}V_{\varepsilon,1}(\widetilde{Z}_{t}^{(n)})\Big{]}$ $\leq(1+c_{1}h)\mathbb{E}\Big{[}V_{\varepsilon,1}(\widetilde{Z}_{t_{n}}^{(n)})\Big{]}+c_{1}\cdot\mathbb{E}\bigg{[}\int_{t_{n}}^{t}V_{\varepsilon,1}(\widetilde{Z}_{s}^{(n)})\,\text{d}s\bigg{]}+Ch.$ (3.37)
Now we focus on the case $q>1$. Applying the Itô formula again on $V_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})$ we get
@@
Plugging (3.38) into (3.36) and noting that
$\mathbb{E}\bigg{[}\int_{t_{n}}^{t}2qV_{\varepsilon,q-1}(\widetilde{Z}_{t_{n}}^{(n)})\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\,\text{d}s\bigg{]}=0,$ (3.39)
we have
@@
Using the Young inequality twice as well as the Jensen inequality on $D_{1}$ we get
@@
$\leq 2(q-1)\mathbb{E}\bigg{[}\int_{t_{n}}^{t}\bigg{(}\big{|}-c_{0}+c_{1}Y_{n}\big{|}^{2q}$
$\quad+(2q-1)\Big{|}\int_{t_{n}}^{s}\widetilde{Z}_{r}^{(n)}V_{\varepsilon,q-2}(\widetilde{Z}_{r}^{(n)})\,\text{d}r\cdot\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\Big{|}^{\frac{2q}{2q-1}}\bigg{)}\,\text{d}s\bigg{]}$
$\leq Ch\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})\Big{]}+\mathbb{E}\bigg{[}\int_{t_{n}}^{t}\bigg{(}(2q-2)^{2}\Big{|}h^{-1}\int_{t_{n}}^{s}\widetilde{Z}_{r}^{(n)}V_{\varepsilon,q-2}(\widetilde{Z}_{r}^{(n)})\,\text{d}r\Big{|}^{\frac{q}{q-1}}$
$\quad+(2q-2)\Big{|}h\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\Big{|}^{2q}\bigg{)}\,\text{d}s\bigg{]}$
$\leq Ch\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})\Big{]}+(2q-2)^{2}\int_{t_{n}}^{t}\sup_{t_{n}\leq r\leq s}\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{r}^{(n)})\Big{]}\text{d}s+Ch^{q+1},$
where we used (3.6), (3.27), (3.29), the Hölder inequality and the moment inequality in the last step. Since the estimation of $D_{3}$ is similar and easy, we present it before treating $D_{2}$. By the generalized Itô isometry [8, Chapter 4, B], the Young inequality as well as (3.6), (3.27), (3.29), the Hölder inequality and the moment inequality we deduce that
@@
Employing the Young inequality, (3.27), (3.31) on $D_{21}$ as well as the Hölder inequality and the moment inequality we get
@@
With regard to $D_{22}$, using the Young inequality twice again as well as the Hölder inequality, the moment inequality, (3.6), (3.19) and (3.27) on $D_{22}$ we deduce that
@@
$\leq\tfrac{2q-2}{h}\int_{t_{n}}^{t}\int_{t_{n}}^{s}\mathbb{E}\bigg{[}\big{|}V_{\varepsilon,q}(\widetilde{Z}_{r}^{(n)})\big{|}^{\frac{2q-2}{2q-1}}\cdot\big{|}\widetilde{Z}_{r}^{(n)}-Y_{n}\big{|}^{\frac{2q}{2q-1}}\bigg{]}\,\text{d}r\text{d}s+Ch$
@@
$\quad+\tfrac{(2q-1)(2q-2)}{h}\int_{t_{n}}^{t}\int_{t_{n}}^{s}\mathbb{E}\Bigg{[}\bigg{|}\int_{t_{n}}^{r}\big{(}-c_{0}+c_{1}Y_{n}+f_{h}(Y_{n})\big{)}\,\text{d}r_{1}+\int_{t_{n}}^{r}g_{h}(Y_{n})\,\text{d}W_{r_{1}}\bigg{|}^{2q}\Bigg{]}\,\text{d}r\text{d}s$
$\leq(2q-2)^{2}h\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})\Big{]}+C\int_{t_{n}}^{t}\sup_{t_{n}\leq r\leq s}\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{r}^{(n)})\Big{]}\text{d}s+Ch.$
@@
$\mathbb{E}\Big{[}\big{(}\varepsilon+|Y_{n+1}-c_{-1}hY_{n+1}^{-1}|^{2}\big{)}^{q}\Big{]}$ $=\mathbb{E}\Big{[}\big{(}1+c_{-1}+2c_{-1}h+|Y_{n+1}|^{2}-2c_{-1}h+|Y_{n+1}|^{-2}\big{)}^{q}\Big{]}$ (3.49)
@@
On the other hand,
@@
Thus we conclude that
@@
Let Assumptions 3.2, 3.3, 3.4 hold. Then it holds that
@@ 4 Mean-square convergence rate of the proposed scheme
In this section we present the analysis of the mean-square convergence rate for the proposed scheme on the mesh grid. Two cases are considered including the non-critical case $\kappa+1>2\rho$ and the general critical case $\kappa+1=2\rho$. For both cases, we successfully identify the desired mean-square convergence rate of order one-half for the considered scheme. The main convergence result is formulated as follows.
@@
Let $\{X_{t}\}_{t\in[0,T]}$ and $\{Y_{n}\}_{0\leq n\leq N}$ be defined by (2.2) and (3.1), respectively. Let Assumptions 3.2, 3.3 and 3.4 hold, with additionally one of the following conditions standing:
@@
where $\delta>0$ is some positive constant. By noting that
$\big{\langle}e_{n+1},X_{t_{n+1}}^{-1}-Y_{n+1}^{-1}\big{\rangle}=\int_{0}^{1}-\Big{(}Y_{n+1}+\theta(X_{t_{n+1}}-Y_{n+1})\Big{)}^{-2}\text{d}\theta\cdot(X_{t_{n+1}}-Y_{n+1})^{2}\leq 0$ (4.5)
and that
$\mathbb{E}\big{[}\langle e_{n},(g(X_{t_{n}})-g(Y_{n}))\Delta W_{n}\rangle\big{]}=0,\,\mathbb{E}\big{[}\langle e_{n},R_{n+1}^{(2)}\rangle\big{]}$ $=0,$
$\mathbb{E}\big{[}\big{\langle}f(X_{t_{n}})-f(Y_{n}),(g(X_{t_{n}})-g(Y_{n}))\Delta W_{n}\big{\rangle}\big{]}$ $=0,$
we thus take expectations on both sides of (4.4) and do some rearrangements to arrive at
@@
In view of this, we claim that
$(1+\delta)h\mathbb{E}\big{[}\big{|}g(X_{t_{n}})-g(Y_{n})\big{|}^{2}\big{]}+2h\mathbb{E}\big{[}\langle e_{n},f(X_{t_{n}})-f(Y_{n})\rangle\big{]}\leq\tilde{L}(X_{t_{n}}-Y_{n})^{2}.$ (4.8)
Indeed, for the case $\kappa+1>2\rho$, the assertion is obvious by noting $\kappa-1>2\rho-2$ and $02c_{2}\kappa>$ 0. For the critical case $\kappa+1=2\rho$, the assumption $\tfrac{c_{2}}{c_{3}^{2}}>2\kappa-\tfrac{3}{2}$ implies $\tfrac{c_{2}}{c_{3}^{2}}>\kappa/2$ and thus $\tfrac{2c_{2}\kappa}{c_{3}^{2}\rho^{2}}-1>0$. As a result, one can choose $10<\delta\leq\tfrac{2c_{2}\kappa}{c_{3}^{2}\rho^{2}}-$ 1 such that $(1+\delta)c_{3}^{2}\rho^{2}-2c_{2}\kappa\leq 0$, which ensures that (4.8) still remains valid. Taking (4.8) into account, one can deduce from (4.6) that

### v2

@@ 3 The explicit positivity-preserving Euler-type scheme
$\mathbb{E}\Big{[}V_{\varepsilon,1}(\widetilde{Z}_{t}^{(n)})\Big{]}$ $\leq(1+c_{1}h)\mathbb{E}\Big{[}V_{\varepsilon,1}(\widetilde{Z}_{t_{n}}^{(n)})\Big{]}+{\color[rgb]{0,0,0}(2c_{0}+c_{1})}\cdot\mathbb{E}\bigg{[}\int_{t_{n}}^{t}V_{\varepsilon,1}(\widetilde{Z}_{s}^{(n)})\,\text{d}s\bigg{]}+Ch.$ (3.38)
Now, we focus on the case $q>1$. Applying the Itô formula again on $V_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)}),$ we get
@@
Similar to (3.35), we note that
$\mathbb{E}\bigg{[}\int_{t_{n}}^{t}2qV_{\varepsilon,q-1}(\widetilde{Z}_{t_{n}}^{(n)})\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\,\text{d}s\bigg{]}$ $=2q\int_{t_{n}}^{t}\mathbb{E}\bigg{[}\int_{t_{n}}^{s}V_{\varepsilon,q-1}(\widetilde{Z}_{t_{n}}^{(n)})f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\bigg{]}\,\text{d}s$ (3.40)
$=0.$
Taking this into account and plugging (3.39) into (3.37), we have
@@
Utilizing the Young inequality of the form
$a\cdot b\leq\tfrac{1}{2q}a^{2q}+\tfrac{2q-1}{2q}b^{\frac{2q}{2q-1}}=\tfrac{1}{2q}\Big{(}a^{2q}+(2q-1)b^{\frac{2q}{2q-1}}\Big{)},$
we first derive
@@
$\leq 2(q-1)\mathbb{E}\bigg{[}\int_{t_{n}}^{t}\big{|}-c_{0}+c_{1}Y_{n}\big{|}^{2q}\,\text{d}s\bigg{]}$
$\quad+2(q-1)(2q-1)\mathbb{E}\bigg{[}\int_{t_{n}}^{t}\Big{|}\int_{t_{n}}^{s}\widetilde{Z}_{r}^{(n)}V_{\varepsilon,q-2}(\widetilde{Z}_{r}^{(n)})\,\text{d}r\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\Big{|}^{\frac{2q}{2q-1}}\,\text{d}s\bigg{]}.$
Further, applying the Young inequality of another form
$a\cdot b\leq\tfrac{2q-2}{2q-1}a^{\frac{2q-1}{2q-2}}+\tfrac{1}{2q-1}b^{2q-1}=\tfrac{1}{2q-1}\big{(}(2q-2)a^{\frac{2q-1}{2q-2}}+b^{2q-1}\big{)},$
we get
$D_{1}$ $\leq 2(q-1)\mathbb{E}\bigg{[}\int_{t_{n}}^{t}\big{|}-c_{0}+c_{1}Y_{n}\big{|}^{2q}\,\text{d}s\bigg{]}$ (3.43)
$\quad+2(q-1)(2q-2)\mathbb{E}\bigg{[}\int_{t_{n}}^{t}\Big{|}h^{-1}\int_{t_{n}}^{s}\widetilde{Z}_{r}^{(n)}V_{\varepsilon,q-2}(\widetilde{Z}_{r}^{(n)})\,\text{d}r\Big{|}^{\frac{2q}{2q-2}}\,\text{d}s\bigg{]}$
$\quad+2(q-1)\mathbb{E}\bigg{[}\int_{t_{n}}^{t}\Big{|}h\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\Big{|}^{2q}\,\text{d}s\bigg{]}.$
For the first term, we can directly deduce that
$|-c_{0}+c_{1}Y_{n}|^{2q}\leq C\big{(}\varepsilon+|Y_{n}|^{2}\big{)}^{q}=CV_{\varepsilon,q}(Y_{n})=CV_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)}).$
For the second term, utilizing (3.27), (3.29) and the Hölder inequality, we get
$\mathbb{E}\bigg{[}\int_{t_{n}}^{t}\Big{|}h^{-1}\int_{t_{n}}^{s}\widetilde{Z}_{r}^{(n)}V_{\varepsilon,q-2}(\widetilde{Z}_{r}^{(n)})\,\text{d}r\Big{|}^{\frac{2q}{2q-2}}\,\text{d}s\bigg{]}$
$\leq\int_{t_{n}}^{t}\mathbb{E}\bigg{[}\Big{|}h^{-1}\int_{t_{n}}^{s}V_{\varepsilon,q-1}(\widetilde{Z}_{r}^{(n)})\,\text{d}r\Big{|}^{\frac{2q}{2q-2}}\bigg{]}\,\text{d}s$
$\leq\int_{t_{n}}^{t}h^{\frac{q}{q-1}-1}\int_{t_{n}}^{s}\mathbb{E}\Big{[}\big{|}h^{-1}V_{\varepsilon,q-1}(\widetilde{Z}_{r}^{(n)})\big{|}^{\frac{q}{q-1}}\Big{]}\text{d}r\text{d}s$
$\leq\int_{t_{n}}^{t}\sup_{t_{n}\leq r\leq s}\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{r}^{(n)})\Big{]}\text{d}s.$
For the third term, the moment inequality [15, Theorem 7.1] together with (3.6) directly infers that
$\mathbb{E}\bigg{[}\int_{t_{n}}^{t}\Big{|}h\int_{t_{n}}^{s}f_{h}(Y_{n})g_{h}(Y_{n})\,\text{d}W_{r}\Big{|}^{2q}\,\text{d}s\bigg{]}\leq Ch^{q+1}.$
Thus, in summary, we have
$D_{1}\leq Ch\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})\Big{]}+(2q-2)^{2}\int_{t_{n}}^{t}\sup_{t_{n}\leq r\leq s}\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{r}^{(n)})\Big{]}\text{d}s+Ch^{q+1}.$ (3.44)
Since the estimation of $D_{3}$ is similar and easy, we present it before treating $D_{2}$. By the generalized Itô isometry [8, Chapter 4, B], the Young inequality as well as (3.6), (3.27), (3.29), the Hölder inequality, and the moment inequality [15, Theorem 7.1], we deduce that
@@
Employing the Young inequality, (3.27), (3.31) on $D_{21}$ as well as the Hölder inequality and the moment inequality [15, Theorem 7.1], we get
@@
With regard to $D_{22}$, using the Young inequality twice again as well as the Hölder inequality, the moment inequality, (3.6), (3.19), and (3.27) on $D_{22}$, we deduce that
@@
$\leq\tfrac{(2q-2){\color[rgb]{0,0,0}(2q-1)}}{h}\int_{t_{n}}^{t}\int_{t_{n}}^{s}\mathbb{E}\bigg{[}\big{|}V_{\varepsilon,q}(\widetilde{Z}_{r}^{(n)})\big{|}^{\frac{2q-2}{2q-1}}\cdot\big{|}\widetilde{Z}_{r}^{(n)}-Y_{n}\big{|}^{\frac{2q}{2q-1}}\bigg{]}\,\text{d}r\text{d}s+Ch$
@@
$\quad+\tfrac{{\color[rgb]{0,0,0}2q-2}}{h}\int_{t_{n}}^{t}\int_{t_{n}}^{s}\mathbb{E}\Bigg{[}\bigg{|}\int_{t_{n}}^{r}\big{(}-c_{0}+c_{1}Y_{n}+f_{h}(Y_{n})\big{)}\,\text{d}r_{1}+\int_{t_{n}}^{r}g_{h}(Y_{n})\,\text{d}W_{r_{1}}\bigg{|}^{2q}\Bigg{]}\,\text{d}r\text{d}s.$
By the Jensen inequality, the Hölder inequality, the moment inequality [15, Theorem 7.1], and the fact that
$\big{|}-c_{0}+c_{1}Y_{n}\big{|}^{2q}\leq C\cdot\big{(}\varepsilon+|Y_{n}|^{2}\big{)}^{q}=C\cdot\big{(}\varepsilon+|\widetilde{Z}_{t_{n}}^{(n)}|^{2}\big{)}^{q}=C\cdot V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)}),$
one can bound the integrand in last term of (3.48) as follows:
$\mathbb{E}$ $\Bigg{[}\bigg{|}\int_{t_{n}}^{r}\big{(}-c_{0}+c_{1}Y_{n}+f_{h}(Y_{n})\big{)}\,\text{d}r_{1}+\int_{t_{n}}^{r}g_{h}(Y_{n})\,\text{d}W_{r_{1}}\bigg{|}^{2q}\Bigg{]}$
$\leq 2^{2q-1}\mathbb{E}\Bigg{[}\bigg{|}\int_{t_{n}}^{r}\big{(}-c_{0}+c_{1}Y_{n}+f_{h}(Y_{n})\big{)}\,\text{d}r_{1}\bigg{|}^{2q}\Bigg{]}+2^{2q-1}\mathbb{E}\Bigg{[}\bigg{|}\int_{t_{n}}^{r}g_{h}(Y_{n})\,\text{d}W_{r_{1}}\bigg{|}^{2q}\Bigg{]}$
$\leq 2^{2q-1}(r-t_{n})^{2q-1}\int_{t_{n}}^{r}\mathbb{E}\Big{[}\big{|}-c_{0}+c_{1}Y_{n}+f_{h}(Y_{n})\big{|}^{2q}\Big{]}\,\text{d}r_{1}$
$\quad+2^{2q-1}(r-t_{n})^{q-1}\int_{t_{n}}^{r}\mathbb{E}\Big{[}\big{|}g_{h}(Y_{n})\big{|}^{2q}\Big{]}\,\text{d}r_{1}$
$\leq 2^{4q-2}(r-t_{n})^{2q}\mathbb{E}\Big{[}\big{|}-c_{0}+c_{1}Y_{n}\big{|}^{2q}\Big{]}+2^{4q-2}(r-t_{n})^{2q}\mathbb{E}\Big{[}\big{|}f_{h}(Y_{n})\big{|}^{2q}\Big{]}$
$\quad+2^{2q-1}(r-t_{n})^{q}\mathbb{E}\Big{[}\big{|}g_{h}(Y_{n})\big{|}^{2q}\Big{]}$
$\leq 2^{4q-2}(r-t_{n})^{2q}C\cdot\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})\Big{]}+2^{4q-2}(r-t_{n})^{2q}h^{-q}+2^{2q-1}(r-t_{n})^{q}h^{-q}.$
Plugging this into (3.48) shows
${\color[rgb]{0,0,0}D_{22}\leq(2q-2)^{2}\int_{t_{n}}^{t}\sup_{t_{n}\leq r\leq s}\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{r}^{(n)})\Big{]}\text{d}s+Ch^{2q+1}\mathbb{E}\Big{[}V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)})\Big{]}+Ch}.$ (3.49)
@@
$\mathbb{E}\Big{[}\big{(}\varepsilon+|Y_{n+1}-c_{-1}hY_{n+1}^{-1}|^{2}\big{)}^{q}\Big{]}$ $=\mathbb{E}\Big{[}\big{(}1+c_{-1}+2c_{-1}h+|Y_{n+1}|^{2}-2c_{-1}h+{\color[rgb]{0,0,0}c_{-1}^{2}h^{2}}|Y_{n+1}|^{-2}\big{)}^{q}\Big{]}$ (3.53)
@@
On the other hand, by noting
$1+c_{-1}+2c_{-1}h+|Y_{n}|^{2}\leq 1+2h+c_{-1}(1+2h)+(1+2h)|Y_{n}|^{2}=(1+2h)\big{(}1+c_{-1}+|Y_{n}|^{2}\big{)},$
one can deduce
@@
Thus, we conclude that
@@
Let Assumptions 3.2, 3.3, and 3.4 hold. Then it holds that
@@ 4 Mean-square convergence rate of the proposed scheme
In this section, we present the analysis of the mean-square convergence rate for the proposed scheme on the mesh grid. Two cases are considered including the non-critical case $\kappa+1>2\rho$ and the general critical case $\kappa+1=2\rho$. For both cases, we successfully identify the desired mean-square convergence rate of order one-half for the considered scheme. The main convergence result is formulated as follows.
@@
Let $\{X_{t}\}_{t\in[0,T]}$ and $\{Y_{n}\}_{0\leq n\leq N}$ be defined by (2.2) and (3.1), respectively. Let Assumptions 3.2, 3.3, and 3.4 hold, with additionally one of the following conditions standing:
@@
where $\delta>0$ is some positive constant. By noting that $e_{n}$, $f(X_{t_{n}})-f(Y_{n})$, $g(X_{t_{n}})-g(Y_{n})$ are $\mathcal{F}_{t_{n}}$-measurable, one can easily see that
$\mathbb{E}\big{[}\langle e_{n},(g(X_{t_{n}})-g(Y_{n}))\Delta W_{n}\rangle\big{]}$ $=0,$
$\mathbb{E}\big{[}\big{\langle}f(X_{t_{n}})-f(Y_{n}),(g(X_{t_{n}})-g(Y_{n}))\Delta W_{n}\big{\rangle}\big{]}$ $=0.$
Since $e_{n}$ is $\mathcal{F}_{t_{n}}$-measurable, the property of stochastic integrals implies that
$\mathbb{E}\big{[}\langle e_{n},R_{n+1}^{(2)}\rangle\big{]}=\mathbb{E}\bigg{[}\int_{t_{n}}^{t_{n+1}}e_{n}\Big{(}\big{(}g(X_{s})-g(X_{t_{n}})\big{)}+\big{(}g(Y_{n})-g_{h}(Y_{n})\big{)}\Big{)}\,\text{d}W_{s}\bigg{]}=0.$
In addition, we note that
$\big{\langle}e_{n+1},X_{t_{n+1}}^{-1}-Y_{n+1}^{-1}\big{\rangle}=\int_{0}^{1}-\Big{(}Y_{n+1}+\theta(X_{t_{n+1}}-Y_{n+1})\Big{)}^{-2}\text{d}\theta\cdot(X_{t_{n+1}}-Y_{n+1})^{2}\leq 0.$ (4.5)
Therefore, we take expectations on both sides of (4.4) and do some rearrangements to arrive at
@@
Here, we claim that for any $u\geq 0$,
$(1+\delta)c_{3}^{2}\rho^{2}u^{2\rho-2}-2c_{2}\kappa u^{\kappa-1}\leq\tilde{L}.$ (4.8)
Indeed, for the case $\kappa+1>2\rho$, the assertion is obvious by noting $\kappa-1>2\rho-2$ and the fact that the coefficient $-2c_{2}\kappa$ of the higher-order term $u^{\kappa-1}$ is negative. For the critical case $\kappa+1=2\rho$, the assumption $\tfrac{c_{2}}{c_{3}^{2}}>2\kappa-\tfrac{3}{2}$ implies that $\tfrac{c_{2}}{c_{3}^{2}}>\tfrac{\kappa}{2}+\tfrac{3}{2}(\kappa-1)>\tfrac{\kappa}{2}$ and thus $\tfrac{2c_{2}\kappa}{c_{3}^{2}\rho^{2}}-1>\tfrac{\kappa^{2}}{\rho}-1>\rho-1>0$. As a result, one can choose $10<\delta\leq\tfrac{2c_{2}\kappa}{c_{3}^{2}\rho^{2}}-$ 1 such that $(1+\delta)c_{3}^{2}\rho^{2}-2c_{2}\kappa\leq 0$, which ensures that the coefficient of $u^{\kappa-1}=u^{2\rho-2}$ is still negative. A combination of the two cases confirms (4.8). Bearing (4.8) in mind, one can readily derive from (4.7) that
$(1+\delta)h\mathbb{E}\big{[}\big{|}g(X_{t_{n}})-g(Y_{n})\big{|}^{2}\big{]}+2h\mathbb{E}\big{[}\langle e_{n},f(X_{t_{n}})-f(Y_{n})\rangle\big{]}\leq\tilde{L}{\color[rgb]{0,0,0}h\mathbb{E}\big{[}|X_{t_{n}}-Y_{n}|^{2}\big{]}}.$ (4.9)
Taking (4.9) into account, one can deduce from (4.6) that
$\mathbb{E}\big{[}|e_{n+1}|^{2}\big{]}$ $\leq(1+(1+3c_{1})h+(c_{1}^{2}+2c_{1})h^{2}+c_{1}^{2}h^{3}{\color[rgb]{0,0,0}+\tilde{L}})\mathbb{E}\big{[}|e_{n}|^{2}\big{]}+(2+c_{1}h)h^{2}\mathbb{E}\big{[}|f(X_{t_{n}})-f(Y_{n})|^{2}\big{]}$ (4.10)